OpenStudy (hatahateo):

Help, I'll fan and medal What are the explicit equation and domain for an arithmetic sequence with a first term of 5 and a second term of 3?

OpenStudy (hatahateo):

Answers: an = 5 − 3(n − 1); all integers where n ≥ 1 an = 5 − 3(n − 1); all integers where n ≥ 0 an = 5 − 2(n − 1); all integers where n ≥ 0 an = 5 − 2(n − 1); all integers where n ≥ 1

OpenStudy (hatahateo):

@mathmate

OpenStudy (hatahateo):

@agent0smith

OpenStudy (mathmate):

So we're given that the arithmetic sequence is {5,3,1,-1,-3,-5,...} Can you identify the common difference? This is important in finding the domain.

OpenStudy (hatahateo):

Umm, that those are negative? @mathmate

OpenStudy (hatahateo):

I'm confused with this question

OpenStudy (mathmate):

Common difference can be a negative number, which means that the sequence becomes progressively smaller. This question essentially asks you to find the general term (the \(n^{th}\) term of the sequence. If you are not familiar with how to find the general term of an arithmetic sequence, the following link gives a very easy-to-read explanation of how you can do that, with examples: https://www.mathsisfun.com/algebra/sequences-sums-arithmetic.html or read more here: http://www.mathcentre.ac.uk/resources/uploaded/mc-ty-apgp-2009-1.pdf